OpenStudy (ledah):

solve by substitution x + 3y = -4 4x + 5y = 5

OpenStudy (amistre64):

what would you like to substitute?

OpenStudy (cwrw238):

x + 3y = -4 ....................(1) 4x + 5y = 5.....................(2) from equation (1) x= -4-3y substitute for x in (2) 4(-4-3y) + 5y = 5 -16 - 12y + 5y = 5 -7y = 21 y = 21/-7 = -3 now plug y = -3 into equation (1) to get x

OpenStudy (amistre64):

or since we have an equation for x in terms of y; just use that one either-or